Solve the following system of equations for the unknown variables. a + b = 7
,a + c = 12,b + c = 15

Answer by ankor@dixie-net.com(22740)

a + b = 7,
a + c = 12,
b + c = 15
:
a + c = 12
a = (12-c)
and
b + c = 15
b = (15-c)
:
The 1st equation
a + b = 7
Replace a and b
(12-c) + (15-c) = 7
-2c + 27 = 7
-2c = 7 - 27
-2c = - 20
c = +10
then
a + 10 = 12
a = 2
and
b + 10 = 15
b = 5
:
You can check these solution in the original equations